# Generalization of terms via universal algebra

### v1

@@ Abstract.
We provide a new foundational approach to the generalization of terms up to equational theories. We interpret generalization problems in a universal-algebraic setting making a key use of projective and exact algebras in the variety associated to the considered equational theory. We prove that the generality poset of a problem and its type (i.e., the cardinality of a complete set of least general solutions) can be studied in this algebraic setting. Moreover, we identify a class of varieties where the study of the generality poset can be fully reduced to the study of the congruence lattice of the 1-generated free algebra. We apply our results to varieties of algebras and to (algebraizable) logics. In particular we obtain several examples of unitary type: abelian groups; commutative monoids and commutative semigroups; all varieties whose 1-generated free algebra is trivial, e.g., lattices, semilattices, varieties without constants whose operations are idempotent; Boolean algebras, Kleene algebras, and Gödel algebras, which are the equivalent algebraic semantics of, respectively, classical, 3-valued Kleene, and Gödel-Dummett logic.
@@ 1. Introduction
A term $s$ is a generalization of a term $t$ if $t$ can be obtained from $s$ through a variable substitution. The problem of identifying common generalizations for two or more terms has been the focus of substantial research, initiated in a series of papers by Plotkin [44], Popplestone [45], and Reynolds [49], all collected in the same volume published in 1970. The objective of these initial papers was to formalize an abstraction of the process of inductive reasoning. As a terminological aside, this abstraction is named “inductive generalization” in [44], whilst more recent literature simply speaks of “generalization processes” (see e.g. [16]). Here we conform to this latter usage. The main aim in this context is to find the solutions, i.e., the generalizing terms, that are as close as possible to the initial terms that define the problem. The cardinality of this set of “best” solutions is encoded in what is called the generalization type.
@@
The present paper provides a novel foundational approach to equational generalization, i.e., where terms are understood to be equivalent up to an equational theory. Our approach is based on universal algebra, which is a most natural environment to handle equational theories from the side of their classes of models, i.e., varieties. The extension to generalization up to equational theories has been considered by several authors in theoretical computer science [5, 11, 46, 8]. Most relevant results so far have been obtained with ad hoc techniques (see, e.g., results on semirings [15] and idempotent operations [16]), and there is a growing interest in a general, foundational approach (see the recent survey [17]).
@@
In the literature, the collection of methods and techniques developed to compute solutions to a generalization problem often goes under the name of anti-unification. This terminology suggests a connection between generalization and the arguably better-known unification problems, where one seeks common instantiations to pairs of given terms. Our approach is indeed inspired by Ghilardi’s algebraic setting for the study of equational unification problems [30]. This has been widely used in the literature since it allows the use of (universal) algebraic techniques [1, 2, 26, 33, 52] and duality theoretic methods employing Priestley-like [10, 13, 12, 32] and geometrical [29, 40, 53] dualities. As we shall show also in the present paper, an advantage given by the universal algebraic perspective is that it brings to light the invariants of both the problems and the solutions: translating a problem to the algebraic setting allows to identify which problems share the same solutions, and which solutions are essentially equivalent up to the considered equational theory.
@@
With our methods we are able to tackle a variety of examples, coming both from algebra and algebraizable logics [9, 28]. Particularly, we show that problems in the following varieties have unitary type: abelian groups, commutative semigroups and monoids (Example 4.26); all varieties whose 1-generated free algebra is trivial, e.g., lattices, semilattices, varieties without constants whose operations are idempotent (Example 4.22); Boolean algebras, Kleene algebras, and Gödel algebras, which are the equivalent algebraic semantics of, respectively, classical, 3-valued Kleene, and Gödel-Dummett logic (Section 5).
@@ 2.1. Universal algebraic preliminaries: projectivity and exactness
$\scriptstyle{j\circ i={\rm id}_{{\textbf{P}}}}$$\scriptstyle{i}$${{\textbf{F}}_{\mathsf{V}}(Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j}$
@@
The relation between congruences, homomorphic images, and quotients of algebras is a basic tool from universal algebra; the main facts are expressed in the so-called Isomorphism Theorems [42, §4.2]. We recall the First Isomorphism Theorem (often called just Homomorphism Theorem) and Second Isomorphism Theorem here for the reader’s convenience, since they will be widely used in proofs.
@@ 2.2. E-generalization problems
(1) $s\preceq u\mbox{, iff there exists a substitution }\sigma\mbox{ such that }\sigma(u)=s.$
@@
Given a poset $(P,\leq)$, we call a minimal (or maximal) complete set a subset $M\subseteq P$ such that:
@@
Given a variety $\mathsf{V}$, its symbolic e-generalization type is the worst possible type occurring among all its e-generalization problems, the best-to-worst order being: unitary $>$ finitary $>$ infinitary $>$ nullary.
@@ 3. An algebraic presentation of e-generalization
In this section we present a novel universal-algebraic approach to e-generalization problems. The starting intuition is that, given a symbolic problem $t_{1},\ldots,t_{m}$, each term defines a 1-generated subalgebra of some free algebra, i.e., a 1-generated exact algebra ${\textbf{E}}(t_{k})$. As we hinted at in Remark 2.8, it is necessary for the study of the generalization order to fix the specific term $t_{k}$, for each $k=1,\ldots,m$, and not just the exact algebra it generates (which could have different generators). The key idea that makes our translation works is to see the terms $t_{1},\ldots,t_{m}$ as a single element $(t_{1},\ldots,t_{m})$ of the direct product of the exact algebras ${\textbf{E}}(t_{k})$; it will be convenient to represent this tuple as the image of a fresh variable $z$ via some homomorphism $h$ on the 1-generated algebra in the associated variety, ${\bf F}_{\mathsf{V}}(z)$. Then, each term $t_{k}$ can be recovered as $p_{k}\circ h(z)$, where $p_{k}:\prod_{k=1}^{m}{\textbf{E}}(t_{k})\to{\textbf{E}}(t_{k})$ is the $k$-th projection.
The next definition translates this intuition in algebraic terms.
@@
${{\bf F}_{\mathsf{V}}(z)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$$\scriptstyle{h}$${\displaystyle\prod_{k=1}^{m}{\textbf{E}}_{k}}$$\scriptstyle{f}$
@@
(2) $g\sqsubseteq g^{\prime}\mbox{ if and only if there exists }f:{\textbf{B}}^{\prime}\to{\textbf{B}}\mbox{ such that }f\circ g^{\prime}=g,$
@@
[…]
@@
$l:=(l_{1}\circ p_{1},\ldots,l_{m}\circ p_{m})\;$ ::: $\;\prod_{k=1}^{m}{\textbf{E}}_{k}\to\prod_{k=1}^{m}{\textbf{E}}(t_{k}),$
$l^{-1}:=(l_{1}^{-1}\circ p_{1},\ldots,l_{m}^{-1}\circ p_{m})\;$ ::: $\;\prod_{k=1}^{m}{\textbf{E}}(t_{k})\to\prod_{k=1}^{m}{\textbf{E}}_{k}$
@@
The reader familiar with the algebraic study of unification problems developed by Ghilardi in [30] will appreciate that our results have a similar flavor, but also some significant differences. Importantly, the study of unification problems and their type only involves homomorphisms between finitely presented and projective algebras in a variety; since these are all categorical notions, it directly follows that the unification type is preserved under categorical equivalence. With the approach outlined above, we do not get this same result, since we make a key use of 1-generated (free) algebras, and being 1-generated is not generally preserved by categorical equivalences. Nonetheless, we prove that the e-generalization type is preserved under a stronger notion of categorical equivalence, called equivalence in [41], which is quite common among algebraic categories. Let us be more precise. For the unexplained notions of category theory we refer the reader to [39].
@@
All the notions involved in the definition of an algebraic e-generalization problem, its solutions, and the generality order, are preserved by equivalence; therefore we can conclude this section with the following result.
@@ 4.2. Algebraic solutions via congruences
Let $\mathsf{V}$ be a variety such that every 1-generated exact algebra is projective. Then given any e-generalization problem $h$, $\mathscr{G}(h)=\{\theta\in\operatorname{Con}({\textbf{{\bf F}}}_{\mathsf{V}}(z)):\theta\subseteq\ker(h),\theta\mbox{ projective in }\mathsf{V}\}$.
@@
Moreover, many varieties coming from logic are 1EP; for instance, Heyting algebras (since all exact Heyting algebras are projective, [31, Remark pp. 867-868]), and as we shall see in the last section, Boolean, Gödel, and Kleene algebras. However, one can construct examples of varieties without the 1EP property; the following is due to Tommaso Moraschini (private communication).
@@
Therefore $\mathsf{N}_{3}$ is not 1EP. It is interesting to further notice that, however, S is weakly projective in $\mathsf{N}_{3}$ [^1] We note that we were not able to find in the literature explicit examples of algebras that are weakly projective but not projective, hence this seemed worth observing. Weakly projective algebras are important in the study of primitive quasivarieties (see e.g., [34, Prop. 5.1.24] or [3, Theorem 4.11])., i.e., whenever it is a homomorphic image of an algebra in the variety it is also (isomorphic to) a subalgebra. This is a consequence of the fact that every algebra in $\mathsf{N}_{3}$ satisfies the identity $3x\approx 4x$ (where $nx$ is $x\oplus\ldots\oplus x$, $n$ times). Indeed, let ${\textbf{A}}\in\mathsf{N}_{3}$ and $h$ a surjective homomorphism $h:{\textbf{A}}\to{\textbf{S}}$; consider $a\in A$ such that $h(a)=2$. If $2a=3a$ then the desired embedding $f:{\textbf{S}}\to{\textbf{A}}$ is defined by mapping $2$ to $a$, and otherwise, since necessarily $3a=4a$, $f$ is defined by mapping $2$ to $2a$. $\Box$
@@
${{\bf F}_{\mathsf{V}}(z)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$$\scriptstyle{g}$$\scriptstyle{n_{\ker(g)}}$${\displaystyle\prod_{k=1}^{m}{\textbf{E}}_{k}}$${{\bf F}_{\mathsf{V}}(z)/\ker(g)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{e}$$\scriptstyle{f^{\prime}}$$\scriptstyle{f}$$\scriptstyle{j}$
Figure 1.
Note that, if the variety is 1EP, $n_{\ker(g)}$ is a solution whenever $\ker(g)\in\mathscr{G}(h)$. Indeed, ${\bf F}_{\mathsf{V}}(z)/\ker(g)$ is projective, and by the Second Isomorphism Theorem, there is $f^{\prime}$ such that $f^{\prime}\circ n_{\ker(g)}=h$ which testifies that $n_{\ker(g)}$ is a solution. Now, if there exists $j:{\textbf{P}}\to{\bf F}_{\mathsf{V}}(z)/\ker(g)$ that makes ${\bf F}_{\mathsf{V}}(z)/\ker(g)$ an $(e,j)$-retract of P, then the composition $f^{\prime}\circ j$ testifies that $g$ is a solution as well. This circle of ideas inspires the following definition. Thus, we identify a class of varieties where we can always perform the last step of the above reasoning, i.e., ${\bf F}_{\mathsf{V}}(z)/\ker(g)$ being an $(e,j)$-retract of P for some $j$.

### v2

@@ Abstract.
We provide a new foundational approach to the generalization of terms up to equational theories. We interpret generalization problems in a universal-algebraic setting making a key use of projective and exact algebras in the variety associated to the considered equational theory. We prove that the generality poset of a problem and its type (i.e., the cardinality of a complete set of least general solutions) can be studied in this algebraic setting. Moreover, we identify a class of varieties where the study of the generality poset can be fully reduced to the study of the congruence lattice of the 1-generated free algebra. We apply our results to varieties of algebras and to (algebraizable) logics. In particular we obtain several examples of unitary type: abelian groups; commutative monoids and commutative semigroups; all varieties whose 1-generated free algebra is trivial, e.g., lattices, semilattices, varieties without constants whose operations are idempotent; Boolean algebras, Kleene algebras, and Gödel algebras, which are the equivalent algebraic semantics of, respectively, classical, 3-valued Kleene, and Gödel-Dummett logic. Finally, we prove that the variety of MV-algebras, the equivalent algebraic semantics of Łukasiewicz logic, has nullary type.
@@ 1. Introduction
A term $s$ is a generalization of a term $t$ if $t$ can be obtained from $s$ through a variable substitution. The problem of identifying common generalizations for two or more terms has been the focus of substantial research, initiated in a series of papers by Plotkin [51], Popplestone [52], and Reynolds [56], all collected in the same volume published in 1970. The objective of these initial papers was to formalize an abstraction of the process of inductive reasoning. As a terminological aside, this abstraction is named “inductive generalization” in [51], whilst more recent literature simply speaks of “generalization processes” (see e.g. [19]). Here we conform to this latter usage. The main aim in this context is to find the solutions, i.e., the generalizing terms, that are as close as possible to the initial terms that define the problem. The cardinality of this set of “best” solutions is encoded in what is called the generalization type. In the literature, the collection of methods and techniques developed to compute solutions to a generalization problem often goes under the name of anti-unification. This terminology suggests a connection between generalization and the arguably better-known unification problems, where one seeks common instantiations to pairs of given terms.
@@
The present paper provides a novel foundational approach to equational generalization, i.e., where terms are understood to be equivalent up to an equational theory. Our approach is based on universal algebra, which is a most natural environment to handle equational theories from the side of their classes of models, i.e., varieties.
The extension to generalization up to equational theories has been considered by several authors in theoretical computer science (see [6, 13, 53, 9, 12]. Most relevant results so far have been obtained with ad hoc techniques (see, e.g., results on semirings [18] and idempotent operations [19]), and there is a growing interest in the general foundational aspects of it, as highlighted in the recent survey [20]. We point out that our approach to equational theories is the usual one adopted in universal algebra, and as such it has a different flavor from the one that has been until now considered in the aforementioned literature. In this work, we consider arbitrary equational theories over fixed languages, and terms are completely identified up to equational theories as elements of free algebras; in the literature, one often finds results about languages where one could have any number of operations, but they all satisfy the same equations such as associativity (A), commutativity (C), and their combination (AC). In this setting terms are thought of as being rewritten by means of such equations. This is a methodological difference more than it is a substantial one, but it does shift the presentation and offers a different perspective.
@@
The foundational approach proposed in the present paper is inspired by Ghilardi’s algebraic setting for the study of equational unification problems [34]. This has been widely used in the literature since it allows the use of (universal) algebraic techniques [2, 3, 30, 37, 59] and duality theoretic methods employing Priestley-like [11, 16, 15, 36] and geometrical [33, 45, 60] dualities. As we shall show also in the present paper, an advantage given by the universal algebraic perspective is that it brings to light the invariants of both the problems and the solutions: translating a problem to the algebraic setting allows to identify which problems share the same solutions, and which solutions are essentially equivalent up to the considered equational theory.
@@
With our methods we are able to tackle a variety of examples, coming both from algebra and algebraizable logics [10, 32]. Particularly, we show that problems in the following varieties have unitary type: abelian groups, commutative semigroups and monoids (Example 4.26); all varieties whose 1-generated free algebra is trivial, e.g., lattices, semilattices, varieties without constants whose operations are idempotent (Example 4.22); Boolean algebras, Kleene algebras, and Gödel algebras, which are the equivalent algebraic semantics of, respectively, classical, 3-valued Kleene, and Gödel-Dummett logic (Section 5). Finally, we provide an interesting example of nullary type coming from logic: the variety of MV-algebras, which is the equivalent algebraic semantics of Łukasiewicz infinite-valued logic. This result is obtained using the well-known Marra-Spada duality between finitely presented MV-algebras and a category of rational polyhedra [45].
@@ 2.1. Universal algebraic preliminaries: projectivity and exactness
[…]
@@
The relation between congruences, homomorphic images, and quotients of algebras is a basic tool from universal algebra; the main facts are expressed in the so-called Isomorphism Theorems [47, §4.2]. We recall the First Isomorphism Theorem (often called just Homomorphism Theorem) and Second Isomorphism Theorem here for the reader’s convenience, since they will be widely used in proofs.
@@ 2.2. E-generalization problems
[…]
@@
Given a poset $(P,\leq)$, we call a minimal (resp. maximal) complete set a subset $M\subseteq P$ such that:
@@
Given a variety $\mathsf{V}$, its symbolic e-generalization type is the worst possible type occurring among all its e-generalization problems, the best-to-worst order being: unitary >> finitary >> infinitary >> nullary.
@@ 3. An algebraic presentation of e-generalization
In this section we present a novel universal-algebraic approach to e-generalization problems. The starting intuition is that, given a symbolic problem $t_{1},\ldots,t_{m}$, each term defines a 1-generated subalgebra of some free algebra, i.e., a 1-generated exact algebra ${\textbf{E}}(t_{k})$. As we hinted at in Remark 2.8, it is necessary for the study of the generalization order to fix the specific term $t_{k}$, for each $k=1,\ldots,m$, and not just the exact algebra it generates (which could have different generators). The idea that makes our translation work is to see the terms $t_{1},\ldots,t_{m}$ as a single element $(t_{1},\ldots,t_{m})$ of the direct product of the exact algebras ${\textbf{E}}(t_{k})$; it will be convenient to represent this tuple as the image of a fresh variable $z$ via some homomorphism $h$ on the 1-generated algebra in the associated variety, ${\bf F}_{\mathsf{V}}(z)$. Then, each term $t_{k}$ can be recovered as $p_{k}\circ h(z)$, where $p_{k}:\prod_{k=1}^{m}{\textbf{E}}(t_{k})\to{\textbf{E}}(t_{k})$ is the $k$-th projection.
Hence, the tuple of terms can be substituted by a tuple of surjective homomorphisms $h_{k}:{\bf F}_{\mathsf{V}}(z)\to{\textbf{E}}(t_{k})$, or, equivalently and more compactly, by a homomorphism $h:{\bf F}_{\mathsf{V}}(z)\longrightarrow\prod_{k=1}^{m}{\textbf{E}}(t_{k})$ such that $p_{k}\circ h$ is surjective on ${\textbf{E}}(t_{k})$. The next definition translates this intuition in more abstract algebraic terms.
@@
${{\bf F}_{\mathsf{V}}(z)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{k}\circ h}$$\scriptstyle{h}$${\displaystyle\prod_{k=1}^{m}{\textbf{E}}_{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{k}}$${{\textbf{E}}_{k}}$
@@
Remark 3.2.
We observe that an equivalent presentation of an algebraic problem would be to consider a tuple $\{h_{k}\}_{k=1}^{m}$ where each $h_{k}:{\bf F}_{\mathsf{V}}(z)\to{\textbf{E}}_{k}$ is a surjective homomorphism to a 1-generated exact algebra. With respect to this presentation, a solution for $\{h_{k}\}_{k=1}^{m}$ is a homomorphism $g:{\bf F}_{\mathsf{V}}(z)\longrightarrow{\textbf{P}}$, where P is finitely generated and projective in $\mathsf{V}$, for which there exists a family of homomorphisms $\{f_{k}\}_{k=1}^{m}$ with $f_{k}:{\textbf{P}}\to{\textbf{E}}_{k}$ such that $f_{k}\circ g=h_{k}$. We picked the above more compact presentation, as it will make clearer some of the general results shown in the next section. Nonetheless, it is useful to keep in mind this different but equivalent perspective. We will make use of it in one of the applications, namely the case of Łukasiewicz logic, in Section 5.4.
@@
[…]
@@
$\mathsf{Alg}({\bf t}):z\in{\bf F}_{\mathsf{V}}(z)$ $\longmapsto$ $(t_{1},\ldots,t_{m})\in\prod_{k=1}^{m}{\textbf{E}}(t_{k}),$
$\mathsf{Alg}(s):z\in{\bf F}_{\mathsf{V}}(z)$ $\longmapsto$ $s\in{\bf F}_{\mathsf{V}}(Y).$
@@
$l:=(l_{1}\circ p_{1},\ldots,l_{m}\circ p_{m})\;$ $:$ $\;\prod_{k=1}^{m}{\textbf{E}}_{k}\to\prod_{k=1}^{m}{\textbf{E}}(t_{k}),$
$l^{-1}:=(l_{1}^{-1}\circ p_{1},\ldots,l_{m}^{-1}\circ p_{m})\;$ $:$ $\;\prod_{k=1}^{m}{\textbf{E}}(t_{k})\to\prod_{k=1}^{m}{\textbf{E}}_{k}$
@@
The reader familiar with the algebraic study of unification problems developed by Ghilardi in [34] will appreciate that our results have a similar flavor, but also some significant differences. Importantly, the study of unification problems and their type only involves homomorphisms between finitely presented and projective algebras in a variety; since these are all categorical notions, it directly follows that the unification type is preserved under categorical equivalence. With the approach outlined above, we do not get this same result, since we make a key use of 1-generated (free) algebras, and being 1-generated is not generally preserved by categorical equivalences. Nonetheless, the e-generalization type is preserved under a stronger notion of categorical equivalence, called equivalence in [46], which is quite common among algebraic categories. Let us be more precise. For the unexplained notions of category theory we refer the reader to [44].
@@
All the notions involved in the definition of an algebraic e-generalization problem, its solutions, and the generality order, are preserved by equivalence; therefore we have the following result.
@@
Remark 3.10.
We conclude this section with the following observation. One could have restricted the algebraic solutions in Definition 3.1 to be homomorphisms among free algebras, $g:{\textbf{F}}_{\mathsf{V}}(z)\to{\textbf{F}}_{\mathsf{V}}(Y)$, for some finite set $Y$. This choice would yield equivalent results with respect to the study of the generality poset of solutions. Indeed, for the non-trivial side, consider a solution $g:{\textbf{F}}_{\mathsf{V}}(z)\to{\textbf{P}}$ for P finitely generated and projective. Then P is an $(i,j)$-retract of some finitely generated free algebra ${\textbf{F}}_{\mathsf{V}}(Y)$. We claim that the composition $i\circ g:{\textbf{F}}_{\mathsf{V}}(z)\to{\textbf{F}}_{\mathsf{V}}(Y)$ is equally general to $g$. Indeed, while by definition of the order it follows directly that $i\circ g\sqsubseteq g$, notice that $g=j\circ(i\circ g)$ and hence $g\sqsubseteq i\circ g$.
However, the presentation involving projective algebras has several advantages. First, it allows one to study generalization problems in varieties where one does not have a handy description of free algebras, but instead has a clear characterization of projective ones. For instance, in some locally finite varieties the finitely generated projective algebras are all the finite algebras, but there is no general description of free algebras (see the example of hoops in Section 5.2); in Section 5.3, we use a dualized representation of projective algebras to characterize the type of Kleene algebras.
Finally, another important observation is that projective algebras are preserved by categorical equivalences among algebraic categories, while free algebras are not. Therefore, we prefer to use the presentation of the solutions with projective algebras; the reader may keep in mind, however, that a presentation with free algebras alone would be equivalent.
@@ 4.2. Algebraic solutions via congruences
Let $\mathsf{V}$ be a variety such that every 1-generated exact algebra is projective. Then given any e-generalization problem $h$,
$\mathscr{G}(h)=\{\theta\in\operatorname{Con}({\textbf{{\bf F}}}_{\mathsf{V}}(z)):\theta\subseteq\ker(h),\theta\mbox{ projective in }\mathsf{V}\}.$
@@
Moreover, many varieties coming from logic are 1EP; for instance, Heyting algebras (since all exact Heyting algebras are projective, [35, Remark pp. 867-868]), and as we shall see in the last section, Boolean, Gödel, and Kleene algebras. However, one can construct examples of varieties without the 1EP property; the following is due to Tommaso Moraschini (private communication).
@@
Therefore $\mathsf{N}_{3}$ is not 1EP. It is interesting to further notice that, however, S is weakly projective in $\mathsf{N}_{3}$ [^1] We note that we were not able to find in the literature explicit examples of algebras that are weakly projective but not projective, hence this seemed worth observing. Weakly projective algebras are important in the study of primitive quasivarieties (see e.g., [38, Prop. 5.1.24] or [4, Theorem 4.11])., i.e., whenever it is a homomorphic image of an algebra in the variety it is also (isomorphic to) a subalgebra. This is a consequence of the fact that every algebra in $\mathsf{N}_{3}$ satisfies the identity $3x\approx 4x$ (where $nx$ is $x\oplus\ldots\oplus x$, $n$ times). Indeed, let ${\textbf{A}}\in\mathsf{N}_{3}$ and $h$ a surjective homomorphism $h:{\textbf{A}}\to{\textbf{S}}$; consider $a\in A$ such that $h(a)=2$. If $2a=3a$ then the desired embedding $f:{\textbf{S}}\to{\textbf{A}}$ is defined by mapping $2$ to $a$, and otherwise, since necessarily $3a=4a$, $f$ is defined by mapping $2$ to $2a$. $\Box$
@@
[…]
Figure 1. Solutions in 1EP varieties.
Note that, if the variety is 1EP, $n_{\ker(g)}$ is a solution whenever $\ker(g)\in\mathscr{G}(h)$. Indeed, ${\bf F}_{\mathsf{V}}(z)/\ker(g)$ is projective, and by the Second Isomorphism Theorem, there is $f^{\prime}$ such that $f^{\prime}\circ n_{\ker(g)}=h$ which testifies that $n_{\ker(g)}$ is a solution. Now, if there exists $j:{\textbf{P}}\to{\bf F}_{\mathsf{V}}(z)/\ker(g)$ that makes ${\bf F}_{\mathsf{V}}(z)/\ker(g)$ an $(e,j)$-retract of P, then the composition $f^{\prime}\circ j$ testifies that $g$ is a solution as well. Thus, we identify a class of varieties where we can always perform the last step of the above reasoning, i.e., ${\bf F}_{\mathsf{V}}(z)/\ker(g)$ being an $(e,j)$-retract of P for some $j$.